OpenStudy (anonymous):

Does there exist a function which is continuous everywhere but not differentiable at exactly two points ?

OpenStudy (merchandize):

Yes it can exist. If you have a piecewise function with three parts, the endpoints can be continuous but not differentiable. For instance, say that you have a linear function that goes from negative infinity to negative 2 and stops. Then for the second part of the piecewise function, from negative 2 to positive 2 you would have a quadratic function. And from positive 2 to infinity, you could have another linear function. At x-values of negative and positive 2, the graphs intersect; therefore, they are continuous. However, at these points, cusps could exist, thereby making the function/curve not differentiable at those two points.

OpenStudy (anonymous):

So, if there is a function like \[f(x)=\sqrt{x^{2}-4}\]then \[x \epsilon [-2,2]\]and at -2 and 2 the function is not differentiable.

OpenStudy (anonymous):

@merchandize ?

OpenStudy (anonymous):

Yes but it is not continuous everywhere.

OpenStudy (anonymous):

Think of setting up a piecewise function as merchandize suggested using the absolute value function.

OpenStudy (merchandize):

yes.

OpenStudy (anonymous):

what is a piecewise function actually ?

OpenStudy (anonymous):

Example: \(f(x) = \left\{\begin{array}{ccc}0 & ; & x \leq 0 \\ 1 & ; & x > 0\end{array}\right.\)

OpenStudy (merchandize):

You can create functions that behave differently depending on the input (x) value

OpenStudy (merchandize):

it is same in the form as Alchemista mentioned in the example.

OpenStudy (anonymous):

 then the graph would be this way right ? @Alchemista

OpenStudy (anonymous):

Correct can you think of a way to modify my example using the absolute value function?

OpenStudy (anonymous):

We want to use the absolute value function to create two points where the function is not differentiable.

OpenStudy (anonymous):

I will give you half of what you need. $$f(x) = \left\{\begin{array}{ccc}|x + 1| & ; & x \leq 0 \\ 1 & ; & x > 0\end{array}\right.$$

OpenStudy (anonymous):

Actually that's all you need.

OpenStudy (anonymous):

Graph this function carefully and try to understand why its not differentiable at -1 and 0 $$f(x) = \left\{\begin{array}{ccc}|x + 1| & ; & x \leq 0 \\ 1 & ; & x > 0\end{array}\right.$$

OpenStudy (anonymous):



OpenStudy (anonymous):

these are the points where the graph turns.

OpenStudy (anonymous):

And they are cusps. In other words they are not smooth so the limit of the difference function will not exist.

OpenStudy (anonymous):

oh yes the left and the right limits won't coincide.

OpenStudy (anonymous):

In less rigorous terms you will end up with a limit where the slope is 0 in one direction and nonzero in the other direction.

OpenStudy (anonymous):

At zero that is.

OpenStudy (anonymous):

now i get it :D

OpenStudy (anonymous):

thank you so much @Alchemista and @merchandize :)